OpenStudy (anonymous):

A population of a colony of bacteria increases by 20 percent every 3 minutes. If at 9:00am the colony had a population of 144,000, what was the population of the colony at 8:54am?

OpenStudy (anonymous):

Ok, well.. Simplest way I see to do it: we know that 6 minutes passed between 8:54am to 9:00am Let's find out what the population was 3 minutes earlier at 8:57am. We know it grew up to 144,000 by those 3 minutes. This increase is 20% of what it was at 8:57am.. so: \[ 120\% = 144,000 \\ 1\% = \frac{144,000}{120} = 1,200 \\ 100\% = 1\% \cdot 100 = 1,200 \cdot 100 = 120,000 \] So at 8:57am we had population of 120,000. Now same idea 3 minutes earlier: \[ 120\% = 120,000 \\ 1\% = \frac{120,000}{120} = 1,000 \\ 100\% = 1\% \cdot 100 = 1,000 \cdot 100 = 100,000 \] If my calculations are right, the population should be 100,000 at 8:54am

OpenStudy (anonymous):

Pitamar is right, if you calculate it as a check, you start at 100,000 + 20% = 120,000 -> 8:57 then another 3 minutes 120,000 + 20% = 120 000 +24 000 = 144 000

OpenStudy (anonymous):

Ye.. well, that's the 'simple' way doing it, but if they would ask you what was it at 8:03am it's not practical. So we could define a function for this f(t), where 't' is the number of minutes passed from 9:00am \[ f(t) = 144,000 \cdot \bigg( \frac{6}{5} \bigg)^{\Large \frac{t}{3}} \] By the given data the result is known to be true only if 't' is divisible by 3. Then to calculate 8:54 we do: \[ 8:54am - 9:00am = -6_{minutes} \\ f(-6_{minutes}) = 144,000 \cdot \bigg( \frac{6}{5} \bigg)^{\Large \frac{-6}{3}} = \\ = 144,000 \cdot \bigg( \frac{6}{5} \bigg)^{-2} = 144,000 \cdot \bigg( \frac{5}{6} \bigg)^{2} = \\ = 144,000 \cdot \frac{25}{36} = \frac{3,600,000}{36} = 100,000 \]

OpenStudy (anonymous):

thanks pitamar , this formula was just what i was looking for. I did solve it using simple way as u did before.

OpenStudy (anonymous):

actually i did it a bit more simply. If x = original quantity , we know 20 % change occured twice, so x*(1.2)*(1.2) = 144000. So solving for x we get x = 100000

OpenStudy (anonymous):

Ye.. I guess I tend to do it over-detailed.. Well glad i could help